The sum of 4 times joans age and 3 times jim's age is 47. Jim is 1 yea
r less than twice as old as joan. Find each of their ages.

Answer by mananth(16946)

The sum of 4 times joans age and 3 times jim's age is 47. Jim is 1 yea
r less than twice as old as joan. Find each of their ages.
...
Joan x years
Jim 2x-1
..
4 times joan's age = 4x
3 times jim's age = 3(2x-1)
..
4x+3(2x-1)=47
4x+6x-3=47
10x-3=47
add 3 to both sides
10x=50
/10
x=5 Joan's age
Jim's age = 2x-1=2*5-1=9
Joan 5 years
Jim 9 years
..
m.ananth@hotmail.ca